wite 4 × 9 × 125 index form
Answer by Edwin McCravy(20066)


  4 =  2×2 = 2²
  9 =  3×3 = 3²
125 = 5×25 = 5×5×5 = 5³

So index form is 2²×3²×5³

Edwin